The sum of the first three terms of a geometric progression (GP) is 38; and the fourth term exceeds the first by 19. find the value of the first term and of the common ratio.

Answer by htmentor(1343)

The nth term of a GP is written a_n = a*r^(n-1), where a=the 1st term, r=common ratio
The sum of the 1st 3 terms = 38 = a + ar + ar^2 = a(1+r+r^2)
The 4th term exceeds the 1st term by 19:
ar^3 = a + 19 -> a(r^3-1) = 19
r^3-1 can be factored as (r-1)(r^2+r+1)
So we have the following two equations:
a(1+r+r^2) = 38
a(r-1)(1+r+r^2) = 19
Divide the 2nd equation by the 1st:
r-1 = 1/2
r = 3/2
The first term = a = 38/(1+3/2+(3/2)^2)) -> a = 38/(19/4) = 8